solve the system by graphing
2x-2y=-2
2x+3y=23

Answer by mananth(16949)

2x-2y=-2
2x+3y=23
equation1 -2 = -2
when x= 0, y=1(0,1)
when y = 0,x= -1(-1,0)
Equation 2
when x= 0, y=7.67(0,7.67)
when y =0,x=11.5(11.5,0)
Plot the points and draw the line

The solution is (4,5)